5*(29). It has been a long time since I have done this type of problem.
Answer by AnlytcPhil(1810)

5*(29) =

5*(2*2*2*2*2*2*2*2*2) =

5*(2*2*2*2*2*2*2*4) =

5*(2*2*2*2*2*2*8) =

5*(2*2*2*2*2*16) =

5*(2*2*2*2*32) =

5*(2*2*2*64) =

5*(2*2*128) =

5*(2*256) =

5*(512)

2560

Edwin
AnlytcPhil@aol.com